# Distribution is Unlimited. PA# AFRL-2024-2340

### v1

@@ IV.3 Topological Conjugacy
Let $u$ be observations of an invertible dynamical system. If the ESP holds between $u$ and $x$, then the reservoir is synchronized (in the generalized sense) with said dynamical system. To avoid additional layers of complexity, we will proceed assuming that $u\in U\subset\mathbb{R}^{M}$ are state observations of some system of interest, which evolves according to $f:U\to U$ on the attractor $U$. We refer the reader to Grigoryeva et al. [18] for a thorough treatment on more general types of observation.
Consider the reservoirs discussed in the previous section, with the appropriate constraints such that the ESP holds and is unique. Then, there is a unique homeomorphic map $\Phi:U\to X$, induced by the dynamics of the driven reservoir, that maps the observations to the reservoir space. Corresponding to this $\Phi$, there are some autonomous dynamics $\Gamma:X\to X$ with attractor $X$ in the reservoir space.
(A) ${U}$${U}$${X}$${X}$$\scriptstyle{f}$$\scriptstyle{\Phi}$$\scriptstyle{\Gamma\circ\Phi}$$\scriptstyle{\Phi}$$\scriptstyle{\Gamma}$       (B) ${U}$${U}$${X}$${X}$$\scriptstyle{f}$$\scriptstyle{\Phi}$$\scriptstyle{\Gamma}$$\scriptstyle{\Phi^{-1}\circ\Gamma}$$\scriptstyle{\Phi^{-1}}$
@@
Since $\Phi$ is a homeomorphism, these dynamics $\Gamma=\Phi\circ f\circ\Phi^{-1}$ are topologically conjugate to $f$. The training phase of the reservoir proceeds assuming that $\Phi$ is a homeomorphism, then approximates $\Phi^{-1}$ by an affine map with matrix $W:\mathbb{R}^{N}\to\mathbb{R}^{M}$ and offsets $b\in\mathbb{R}^{M}$. Both $W$ and $b$ are acquired using regularized least squares. As discussed around (23), solving for $W$ and $b$ allows for the definition of $\bar{\sigma}(x_{t})\equiv\sigma(\bar{A}x_{t}+\bar{b})$, which is compatible with the UAP and an attempt to approximate the true conjugate dynamics $\Gamma$. Define the map $\bar{W}(x_{t})=Wx_{t}+b$, not only does $\bar{W}$ approximate $\Phi^{-1}$, but one also gets $\bar{W}\circ\bar{\sigma}$ as an approximation to $\Phi^{-1}\circ\Gamma$ without additional effort.
Despite all this, the map $\Phi$ is only induced by the dynamics when data $u$ is presented to the non-autonomous reservoir. Figure 1(C) holds as an arbitrarily good approximation of Figure 1(B) for the first time-step of the prediction phase. After the first step, the slightest and inevitable deviation of $\bar{W}\circ\bar{\sigma}$ from $\Phi^{-1}\circ\Gamma$ breaks the topological conjugacy. Such errors will compound, resulting in the eventual divergence of the initially conjugated dynamics from the original dynamics; this is distinct from the divergence due to the chaotic dynamics. In the absence of a continuous input of data, this is unavoidable save for the case of the perfect representation of $\Phi^{-1}\circ\Gamma$.
@@ V Conclusions
This work outlines how global nonlinear state-contraction, while driven by inputs, can be uniformly guaranteed for both discrete and continuous time reservoirs. Particularly, a discrete time reservoir with any contracting activation function $\sigma$, any connectivity matrix $A:\min_{i}\norm{A}_{i}<1$, and any $B$ will result in the ESP and unique GS, regardless of the structure of either matrix or the inputs $u$. In continuous time, we show how the widely used Hopfield-like reservoirs with condition $\norm{A}_{2}<s_{N}(C)$ is sufficient for contraction, before showing how lognorms can also provide sufficient conditions for the $1$-lognorm and $\infty$-lognorm cases. All these conditions induce contraction in continuous time, hence the ESP and GS hold. All of the results with matrix norms and lognorms can be further generalized with their weighted operator norms counterparts in the same manner as Buehner and Young [7]. Towards the end of the paper, we discuss how the UAP may appear in reservoir systems and outline the overlap of continuous time reservoirs with NODEs. This suggests that the universal approximation of homeomorphisms and diffeomorphisms, proven for certain classes of NODEs, is likely inherited by continuous time reservoirs. We also discuss and outline how topological conjugacy appears and remark on the surprising robustness of this conjugacy even when errors are accumulated in the prediction phase.
@@ VI Acknowledgments
The authors would like to thank Justin Koo for helpful discussions, suggestions, and encouragement during his time at the Air Force Research Laboratory. This research was supported by the Air Force Office of Scientific Research under FA9550-23RQCOR001.
@@ Appendix A Contracting Activation Function
The contraction mapping property is an incredibly useful property of a function and is repeatedly used in this work. For readability, a function with this property will be described as contracting or referred to as a contraction. Let $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ be a continuously differentiable function that acts component-wise on its argument $v$, i.e. if $v=[v_{1},v_{2}]^{\top}$, then $f(v)=[f(v_{1}),f(v_{2})]^{\top}$ where the scalar- and vector-valued functions are both called $f$. This function is a contraction with respected to an arbitrarily chosen norm when the following inequality holds for $10\leq k<$ 1.
$\norm{f(a)-f(b)}\leq k\norm{a-b}$ (27)
The smallest such scalar $k$, for all $a\neq b$ in $\mathbb{R}^{N}$, is the Lipschitz constant of $f$ and determines the rate at which iterated maps of $f$ converge to a unique fixed point. The value of $k$ is also equal to the magnitude of the supremum slope of $f$ in the one-dimensional setting. For the multidimensional case, the inequality (27) states that a contracting $f$ will always shorten the distance between any two vectors, even if a shift was performed on both vectors. Consider briefly a more general form of (27) by including some matrix $M\in\mathbb{R}^{N\times N}$ and offsets $c\in\mathbb{R}^{N}$ inside the argument. The definition of an induced matrix norm gives the following inequality.
@@
$\leq k\norm{M}\!\cdot\!\norm{a-b}$
In order for the above inequality to also be a contraction, consider an activation function $\sigma$ (such as the ubiquitous tanh function) that satisfies $\sigma(0)=0$, $\sigma^{\prime}(0)=1$, and $\max_{x}\sigma^{\prime}(x)=1$, where the Lipschitz constant of $\sigma$ is $k=1$. Without loss of generality, $k$ can be set to unity and contraction would occur for all vectors $a,b\in\mathbb{R}^{N}$ for $M$ satisfying $10\leq\norm{M}<$ 1. This guarantees that $a=b$ is the unique fixed point.
$\norm{\sigma(Ma+c)-\sigma(Mb+c)}\leq\norm{M}\!\cdot\!\norm{a-b}$ (29)
This allows for (29) to be invoked in (4) to show the exponential stability of certain discrete time reservoir systems. We can also multiply the 2-norm version of (29) by $\norm{a-b}_{2}$ and apply Cauchy-Schwarz to recover the left inequality below. This inequality is used in (9) with constraint $\norm{M}\leq 1$ to show the exponential stability of certain continuous time reservoir systems.
@@ Appendix B Singular Value as Upper and Lower Bounds
Let $C=C^{\top}$ be a symmetric positive definite (SPD) matrix with an eigendecomposition $C=Q^{\top}SQ$, where $S$ is a diagonal matrix and $Q$ is an orthonormal matrix. Being SPD, the eigenvalues of $C$ coincide with the singular values $s_{n}$ that populate the diagonal of $S$, ordered such that $s_{1}\geq s_{2}\geq\cdots\geq s_{N}>0$, .
The quantity of interest $\langle Cz,z\rangle=\norm{z^{\top}Cz}_{2}$ can then be written as $\norm{(Qz)^{\top}S(Qz)}_{2}={\sum_{n=1}^{N}s_{n}[Qz]_{n}^{2}}$, where $[Qz]_{n}$ is the $n^{\text{th}}$ component of the vector $Qz$. Since $s_{n}$ is ordered from largest to smallest, the inner product of interest can be written as a sum with an upper and lower bound.
$s_{N}\sum_{n=1}^{N}[Qz]_{n}^{2}\leq\sum_{n=1}^{N}s_{n}[Qz]_{n}^{2}\leq s_{1}\sum_{n=1}^{N}[Qz]_{n}^{2}$ (31)
With $Q$ being orthonormal, it has no effect on the norm of $z$, i.e. $\sum_{n=1}^{N}[Qz]_{n}^{2}=\norm{Qz}_{2}^{2}=\norm{z}_{2}^{2}$. The sum in both the upper and lower bounds are equal to the norm of $\norm{z}_{2}$. The middle term above is equal to $\langle Cz,z\rangle$ by definition, resulting in the following inequalities.
$s_{N}\norm{z}_{2}^{2}\leq\langle Cz,z\rangle\leq s_{1}\norm{z}_{2}^{2}$ (32)
@@ Appendix C Properties of Logarithmic Norms
$\mu(M)$ $=\frac{d}{dh}\norm{I+hM}\bigg{|}_{h\to 0^{+}}$ (34)
$=\frac{d}{dh}\norm{\exp(hM)}\bigg{|}_{h\to 0^{+}}.$
Unlike the matrix norm, the lognorm is not actually a norm. In this work, we only make use of $\mu_{1}(M)$ and $\mu_{\infty}(M)$, which are the lognorms induced by the matrix $1$-norm and $\infty$-norm respectively. The definitions of $\mu_{1}(M)$ and $\mu_{\infty}(M)$ are the maxima of the following arguments over the columns and rows of an arbitrary square matrix $M$, with elements $m_{ij}$, respectively.
Both $\mu_{1}(M)$ and $\mu_{\infty}(M)$, shown in (17) and (20), are important to guarantee the ESP and GS of continuous time reservoirs in (15). The 2-lognorm $\mu_{2}(M)$ has not been successfully applied in this context, but is provided here for context. Its definition is related to the spectral properties of $M+M^{\top}$. Specifically, it is half the maximal (most positive) eigenvalue of the symmetric matrix $M+M^{\top}$.
$\mu_{2}(M)=\frac{1}{2}\lambda_{max}\left({M+M^{\top}}\right)\$ (35)
@@ Appendix D Properties of Weak Pairing
    sub-additivity in the first argument $\langle\!\langle x_{1}+x_{2},y\rangle\!\rangle\leq\langle\!\langle x_{1},y\rangle\!\rangle+\langle\!\langle x_{2},y\rangle\!\rangle$ for all $x_{1},x_{2},y\in\mathbb{R}^{N}$;
@@
    Lumer’s equality for $M\in\mathbb{R}^{N\times N}$ and $x\in\mathbb{R}^{N}$
    $\mu(M)=\max_{x\neq 0}\frac{\langle\!\langle Mx,x\rangle\!\rangle}{\langle\!\langle x,x\rangle\!\rangle}=\max_{\norm{x}=1}\langle\!\langle Mx,x\rangle\!\rangle;$
@@
For continuous but not differentiable trajectories, the derivative in $(i)$ above can be replaced by the appropriate Dini derivative.
@@ Appendix E Maximum of Activation Function Jacobians
The results of this appendix are attributed to Lemma 2.15 of Bullo [25], but with additional details. To achieve a more useful form of the Jacobian $D\hat{\sigma}(x)$ with respect to $x$, we use the shorthand $\sigma(Ax+Bu)=\hat{\sigma}(x)$ and define $\sigma^{\prime}$ as the derivative of the scalar activation function $\sigma$. We also define “diag” as the operation that places its $N$-vector argument along the diagonal of a $N\times N$ diagonal matrix, i.e. diag$(\begin{bmatrix}v_{1},v_{2}\end{bmatrix}^{\top})=\begin{bmatrix}v_{1}&0\\ 0&v_{2}\end{bmatrix}$, and another shorthand for the resulting diagonal matrix $P=\text{diag}(\sigma^{\prime}(Ax+Bu))$ with only $N$ diagonal elements $p_{jj}$.
@@
$=PA$
For an element-wise activation function $\sigma$ with $\sigma^{\prime}\in(0,1]$, such as the commonly used tanh function, the maximum over all $x$ can be replaced with a maximum over all diagonal elements $p_{jj}$ in the interval $(0,1]$. We first handle the 1-lognorm case.
@@
$=\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}$
The matrix $(A-I\circ A)$ makes use of the entry-wise or Hadamard product, where the elements $(I\circ A)_{ij}=(I)_{ij}(A)_{ij}$. More simply, $(A-I\circ A)$ is just $A$ with its diagonal elements set to zero. Next, we handle the $\infty$-norm case. This case differs from the 1-norm case because the elements of the diagonal matrix $P$ act on the rows of $A$, which allow for $p_{ii}$ to be factored out of the row sum. We define yet another shorthand $q_{i}=a_{ii}+\sum_{j\neq i}\absolutevalue{a_{ij}}$ |.
@@
$=\max\{\mu_{\infty}(A),0\}$
Lastly, we provide some bounds using the equivalent matrix norms. The matrix 1-norm is defined as the maximum of the absolute column sum and the matrix $\infty$-norm is defined as the maximum of the absolute row sum.
$\norm{A}_{1}=\max_{j}\left(\sum_{i}\absolutevalue{a_{ij}}\right)\$ (39)
$\norm{A}_{\infty}=\max_{i}\left(\sum_{j}\absolutevalue{a_{ij}}\right)$ (40)

### v2

@@ IV.3 Topological Conjugacy
Let $u$ be observations of an invertible dynamical system. If the ESP holds between $u$ and $x$, then the reservoir is synchronized (in the generalized sense) with said dynamical system. We will proceed assuming that $u\in U\subset\mathbb{R}^{M}$ are direct state observations of some system of interest, which evolves according to $f:U\to U$ on the attractor $U$. Generally speaking, direct state observations are not available, and there would be an intermediary observation function. Though the importance of observation functions cannot be understated, they are intentionally omitted to avoid additional layers of complexity. The added complexity arises from discussions around generic observables, sparse measurements, and measurement noise. We refer the reader to Grigoryeva et al. [18] for a thorough treatment when observation functions are present.
Consider the reservoirs discussed in the previous section, with the appropriate constraints such that the ESP holds and is unique. Then, there is a unique smooth map $\Phi:U\to X$, induced by the dynamics of the driven reservoir, that maps the observations to the reservoir space. This smooth map almost surely results in an embedding, using Theorem 2.2 of Sauer et al. [28], so its inverse must also be smooth. Corresponding to this $\Phi$, there are some autonomous dynamics $\Gamma:X\to X$ with attractor $X$ in the reservoir space.
(a)                      (b)                     (c)
@@
Since $\Phi$ results in an embedding, it is a homeomorphism. In the continuous time case, this would be a diffeomorphism, but we will leave it as a homeomorphism so we can discuss both discrete and continuous time reservoirs. The dynamics $\Gamma=\Phi\circ f\circ\Phi^{-1}$ are topologically conjugate to $f$. The training phase of the reservoir proceeds under the initial assumption that $\Phi$ is a homeomorphism, then approximates $\Phi^{-1}$ by an affine map with matrix $W:\mathbb{R}^{N}\to\mathbb{R}^{M}$ and offsets $b\in\mathbb{R}^{M}$. Both $W$ and $b$ are acquired using regularized least squares. As discussed around (23), solving for $W$ and $b$ allows for the definition of $\bar{\sigma}(x_{t})\equiv\sigma(\bar{A}x_{t}+\bar{b})$, which is compatible with the UAP and an attempt to approximate the true conjugate dynamics $\Gamma$. Define the map $\bar{W}(x_{t})=Wx_{t}+b$, not only does $\bar{W}$ approximate $\Phi^{-1}$, but one also gets $\bar{W}\circ\bar{\sigma}$ as an approximation to $\Phi^{-1}\circ\Gamma$ without additional effort.
Despite all this, the map $\Phi$ is only induced by the dynamics when data $u$ is presented to the nonautonomous reservoir. Fig.1(c) holds as an arbitrarily good approximation of Fig.1(b) for the first time-step of the prediction phase. After the first step, the slightest and inevitable deviation of $\bar{W}\circ\bar{\sigma}$ from $\Phi^{-1}\circ\Gamma$ breaks the topological conjugacy. Such errors will compound, resulting in the eventual divergence of the initially conjugated dynamics from the original dynamics; this is distinct from the divergence due to the chaotic dynamics. In the absence of a continuous input of data, this is unavoidable save for the case of the perfect representation of $\Phi^{-1}\circ\Gamma$.
@@
IV.4 Digital Twins
One of the most fundamental properties of RC is the necessity of contraction whilst driven by some input signal, as demonstrated throughout this paper. As long as data is continually being presented to the reservoir, the reservoir will maintain synchronization with the source system. When the data stream stops, the readout of the RC is fed back as inputs, allowing short-term predictions of the original data stream.
Interestingly, these features of the reservoir make it well suited for usage as a digital twin (DT). We will use the definition of DT put forth by the Committee on Foundational Research Gaps and Future Directions for Digital Twins [43], reproduced here.
“A digital twin is a set of virtual information constructs that mimics the structure, context, and behavior of natural, engineered, or social system (or system-of-systems), is dynamically updated with data from its physical twin, has a predictive capability, and informs decisions that realize value. The bidirectional interaction between the virtual and physical is central to the digital twin.”
An appropriately trained reservoir is clearly capable of mimicking certain structural aspects of the original system, most famously the invariant measures or climate of the attractor[12]. In the listening phase, the reservoir also inherits the topology of the data due to the homeomorphism that contraction induces. To the extent that there is structure in the topology, the reservoir inherits or mimics that structure as well.
The RC framework is agnostic of the type of system that generates the data, as long as there is some low-dimensional structure in the form of an attractor. It has limited success when applied to text, speech, images, and systems governed by partial differential equations. Other than these limitations, RC has be deployed with considerable success on natural, engineered, and social systems.
The requirements of DTs to be “dynamically updated with data from its physical twin” and to have “a predictive capability” itself has striking similarities to how reservoirs operate, specifically the listening and prediction phase of RC. Though all recurrent networks can be updated with data, the presence of contraction, ESP, and GS in the RC framework is noteworthy. With the reservoir synchronized to the inputs, the topological properties of the data is guaranteed to be expressed in the reservoir, a feature that most recurrent networks do not exhibit.
There is the requirement that the virtual and physical twin have bidirectional interactions. The RC framework only allows for the development of the virtual twin from measurements of the physical twin, but leaves out the interaction on the physical using information of the virtual. RC appears to be agnostic yet entirely compatible to this segment of the DT interaction loop, so it is up to the DT’s user to specify the purpose and how this can be accomplished.
Some of these purposes, based on the strengths of RC, might be to leverage its predictive capabilities to anticipate future behavior. If future behavior is undesirable, one can intervene on the physical twin side – say, by changing system parameters – to avoid or correct for the undesirable behavior. Exactly how this would be accomplished is of course domain and system specific. Nonetheless, such options are open and the application space appears promising.
Lastly, there is no mention of the fidelity of the DT, defined roughly by the amount of physics in the model. In general, lower fidelity models are quicker but inherently less physically informative, while higher fidelity models are the converse. It is natural to assume that the multi-fidelity approach be desirable in DTs, since different levels of fidelity have their respective trade-offs. We are optimistic on the prospects of using RC as one of the lower fidelity components of the DT, but not as a replacements of higher fidelity and physically sound models.
@@ V Conclusions
This paper outlines how global nonlinear state-contraction, while driven by inputs, can be uniformly guaranteed for both discrete and continuous time reservoirs. Particularly, a discrete time reservoir with any contracting activation function $\sigma$, any connectivity matrix $A:\min_{i}\norm{A}_{i}<1$, and any $B$ will result in the ESP and unique GS, regardless of the structure of either matrix or the inputs $u$. In continuous time, we show how the widely used Hopfield-like reservoirs with condition $\norm{A}_{2}<s_{N}(C)$ is sufficient for contraction, before showing how lognorms can also provide sufficient conditions for the $1$-lognorm and $\infty$-lognorm cases. All these conditions induce contraction in continuous time, hence the ESP and GS hold. All of the results with matrix norms and lognorms can be further generalized with their weighted operator norms counterparts in the same manner as Buehner and Young [8]. Towards the end of the paper, we discuss how the UAP may appear in reservoir systems and outline the overlap of continuous time reservoirs with NODEs. This suggests that the universal approximation of homeomorphisms and diffeomorphisms, proven for certain classes of NODEs, is likely inherited by continuous time reservoirs. We discuss and outline how topological conjugacy appears and remark on the surprising robustness of this conjugacy even when errors are accumulated in the prediction phase. Lastly, we lament on how the RC framework is strongly suited for DTs, particularly due to the synchronization.
@@ VI Acknowledgments
The authors thank J. Koo for helpful discussions, suggestions, and encouragement during his time at the Air Force Research Laboratory. This research was supported by the Air Force Office of Scientific Research under FA9550-23RQCOR001 and FA9550-23RQCOR007.
@@ Appendix A Contracting Activation Function
The contraction mapping property is an incredibly useful property of a function and is repeatedly used in this paper. For readability, a function with this property will be described as contracting or referred to as a contraction. Let $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ be a smooth function that acts component-wise on its argument $v$, i.e. if $v=[v_{1},v_{2}]^{\top}$, then $f(v)=[f(v_{1}),f(v_{2})]^{\top}$ where the scalar- and vector-valued functions are both called $f$. This function is a contraction with respect to an arbitrarily chosen norm when the following inequality holds for $10\leq k<$ 1:
$\norm{f(a)-f(b)}\leq k\norm{a-b}.$ (27)
The smallest such scalar $k$, for all $a\neq b$ in $\mathbb{R}^{N}$, is the Lipschitz constant of $f$ and determines the rate at which iterated maps of $f$ converge to a unique fixed point. The value of $k$ is also equal to the magnitude of the supremum slope of $f$ in the one-dimensional setting. For the multidimensional case, the inequality (27) states that a contracting $f$ will always shorten the distance between any two vectors, even if a shift was performed on both vectors. Consider briefly a more general form of (27) by including some matrix $M\in\mathbb{R}^{N\times N}$ and offsets $c\in\mathbb{R}^{N}$ inside the argument. The definition of an induced matrix norm gives the following inequality:
@@
$\leq k\norm{M}\!\cdot\!\norm{a-b}.$
For the above inequality to also be a contraction, consider an activation function $\sigma$ (such as the ubiquitous tanh function) that satisfies $\sigma(0)=0$, $\sigma^{\prime}(0)=1$, and $\max_{x}\sigma^{\prime}(x)=1$, where the Lipschitz constant of $\sigma$ is $k=1$. Without loss of generality, $k$ can be set to unity and contraction would occur for all vectors $a,b\in\mathbb{R}^{N}$ for $M$ satisfying $10\leq\norm{M}<$ 1. This guarantees that $a=b$ is the unique fixed point:
$\norm{\sigma(Ma+c)-\sigma(Mb+c)}\leq\norm{M}\!\cdot\!\norm{a-b}.$ (29)
This allows for (29) to be invoked in (4) to show the exponential stability of certain discrete time reservoir systems. We can also multiply the 2-norm version of (29) by $\norm{a-b}_{2}$ and apply Cauchy-Schwarz to recover the left inequality below. This inequality is used in (9) with constraint $10\leq\norm{M}<$ 1 to show the exponential stability of certain continuous time reservoir systems:
@@ Appendix B Singular Value as Upper and Lower Bounds
Let $C=C^{\top}$ be a SPD matrix with an eigendecomposition $C=Q^{\top}SQ$, where $S$ is a diagonal matrix and $Q$ is an orthonormal matrix. Being SPD, the eigenvalues of $C$ coincide with the singular values $s_{n}$ that populate the diagonal of $S$, ordered such that $s_{1}\geq s_{2}\geq\cdots\geq s_{N}>0$, .
The quantity of interest $\langle Cz,z\rangle=\norm{z^{\top}Cz}_{2}$ can then be written as $\norm{(Qz)^{\top}S(Qz)}_{2}={\sum_{n=1}^{N}s_{n}[Qz]_{n}^{2}}$, where $[Qz]_{n}$ is the $n^{\text{th}}$ component of the vector $Qz$. Since $s_{n}$ is ordered from largest to smallest, the inner product of interest can be written as a sum with an upper and lower bound:
$s_{N}\sum_{n=1}^{N}[Qz]_{n}^{2}\leq\sum_{n=1}^{N}s_{n}[Qz]_{n}^{2}\leq s_{1}\sum_{n=1}^{N}[Qz]_{n}^{2}.$ (31)
With $Q$ being orthonormal, it has no effect on the norm of $z$, i.e. $\sum_{n=1}^{N}[Qz]_{n}^{2}=\norm{Qz}_{2}^{2}=\norm{z}_{2}^{2}$. The sum in both the upper and lower bounds are equal to the norm of $\norm{z}_{2}$. The middle term above is equal to $\langle Cz,z\rangle$ by definition, resulting in the following inequalities:
$s_{N}\norm{z}_{2}^{2}\leq\langle Cz,z\rangle\leq s_{1}\norm{z}_{2}^{2}.$ (32)
@@ Appendix C Properties of Logarithmic Norms
$\mu(A)$ $=\frac{d}{dh}\norm{I+hA}\bigg{|}_{h\to 0^{+}}$ (34)
$=\frac{d}{dh}\norm{\exp(hA)}\bigg{|}_{h\to 0^{+}}.$
Unlike the matrix norm, the lognorm is not actually a norm. In this paper, we only make use of $\mu_{1}(A)$ and $\mu_{\infty}(A)$, which are the lognorms induced by the matrix $1$-norm and $\infty$-norm respectively. The definitions of $\mu_{1}(A)$ and $\mu_{\infty}(A)$ are the maxima of the following arguments over the columns and rows of an arbitrary square matrix $A$, with elements $a_{ij}$, respectively.
Both $\mu_{1}(A)$ and $\mu_{\infty}(A)$, shown in (17) and (20), are important to guarantee the ESP and GS of continuous time reservoirs in (15). The 2-lognorm $\mu_{2}(A)$ has not been successfully applied in this context, but is provided here for context. Its definition is related to the spectral properties of $A+A^{\top}$. Specifically, it is half the maximal (most positive) eigenvalue of the symmetric matrix $A+A^{\top}$:
$\mu_{2}(A)=\frac{1}{2}\lambda_{max}\left({A+A^{\top}}\right).\$ (35)
@@ Appendix D Properties of Weak Pairing
    subadditivity in the first argument $\langle\!\langle x_{1}+x_{2},y\rangle\!\rangle\leq\langle\!\langle x_{1},y\rangle\!\rangle+\langle\!\langle x_{2},y\rangle\!\rangle$ for all $x_{1},x_{2},y\in\mathbb{R}^{N};$
@@
    Lumer’s equality for $A\in\mathbb{R}^{N\times N}$ and $x\in\mathbb{R}^{N}$
    $\mu(A)=\max_{x\neq 0}\frac{\langle\!\langle Ax,x\rangle\!\rangle}{\langle\!\langle x,x\rangle\!\rangle}=\max_{\norm{x}=1}\langle\!\langle Ax,x\rangle\!\rangle;$
@@
For continuous but not differentiable trajectories, the derivative in Definition 2$(i)$ above can be replaced by the appropriate Dini derivative.
@@ Appendix E Maximum of Activation Function Jacobians
The results of this appendix are attributed to Lemma 2.15 of Bullo [26], but with additional details. To achieve a more useful form of the Jacobian $D\hat{\sigma}(x)$ with respect to $x$, we use the shorthand $\sigma(Ax+Bu)=\hat{\sigma}(x)$ and define $\sigma^{\prime}$ as the derivative of the scalar activation function $\sigma$. We also define “diag” as the operation that places its $N$-vector argument along the diagonal of a $N\times N$ diagonal matrix, i.e. diag$(\begin{bmatrix}v_{1},v_{2}\end{bmatrix}^{\top})=\begin{bmatrix}v_{1}&0\\ 0&v_{2}\end{bmatrix}$, and another shorthand for the resulting diagonal matrix $P=\text{diag}(\sigma^{\prime}(Ax+Bu))$ with only $N$ diagonal elements $p_{jj}$:
@@
$=PA.$
For an element-wise activation function $\sigma$ with $\sigma^{\prime}\in(0,1]$, such as the commonly used tanh function, the maximum over all $x$ can be replaced with a maximum over all diagonal elements $p_{jj}$ in the interval $(0,1]$. We first handle the 1-lognorm case:
@@
$=\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}.$
The matrix $(A-I\circ A)$ makes use of the entry-wise or Hadamard product, where the elements $(I\circ A)_{ij}=(I)_{ij}(A)_{ij}$. More simply, $(A-I\circ A)$ is just $A$ with its diagonal elements set to zero. Next, we handle the $\infty$-norm case. This case differs from the 1-norm case because the elements of the diagonal matrix $P$ act on the rows of $A$, which allow for $p_{ii}$ to be factored out of the row sum. We define yet another shorthand $q_{i}=a_{ii}+\sum_{j\neq i}\absolutevalue{a_{ij}}$ |:
@@
$=\max\{\mu_{\infty}(A),0\}.$
Lastly, we provide some bounds using the equivalent matrix norms. The matrix 1-norm is defined as the maximum of the absolute column sum and the matrix $\infty$-norm is defined as the maximum of the absolute row sum:
$\norm{A}_{1}=\max_{j}\left(\sum_{i}\absolutevalue{a_{ij}}\right),\$ (39)
$\norm{A}_{\infty}=\max_{i}\left(\sum_{j}\absolutevalue{a_{ij}}\right).$ (40)